# Regularization of the superposition principle: Potential theory meets Fokker-Planck equations

### 05 Mar 2026 v1

@@ 2. $(T4^{\prime})$
Further, $\mathbf{H_{\sf\lesssim u}}$ follows from [70, Theorem 33] presented above, and 4.6, (i).
@@ 6.1 Fundamentals on right processes and their potential theory
Let $(E,\mathcal{B})$ be a Lusin measurable space, i.e. it is measurably isomorphic to a Borel subset of a compact metric space endowed with the corresponding Borel $\sigma$-algebra. We denote by $(b)p\mathcal{B}$ the set of all numerical, (bounded) positive $\mathcal{B}$-measurable functions on $E$. Throughout, by $\mathcal{U}=(U_{\alpha})_{\alpha>0}$ we denote a resolvent family of (sub-)Markovian kernels on $(E,\mathcal{B})$; see e.g. [13, Subsection 1.1]. If $\beta>0$, we set $\mathcal{U}_{\beta}:=(U_{\beta+\alpha})_{\alpha>0}$.
@@
    Let $(E,\mathcal{B})$ be a Lusin measurable space, and $(f_{n})_{n\geq 1}\subset\mathcal{B}$ be such that it separates the points of $E$. Then the topology on $E$ generated by $(f_{n})_{n\geq 1}$ is a Lusin topology; see e.g. [66, Theorem A2.12] for the similar Radonian case.
@@
Indeed, one can construct such $\tilde{X}$ by considering first a second process equivalent to $X$ which is of function space type, according to [33, Definition 4.2 and Theorem 4.3], and then simply deleting from the path space of this second process the set $\Lambda$ of those paths that are not right-continuous with respect to $\tau$. Such a removal is allowed since $\Lambda$ is universally measurable according to [41, Ch. IV, Theorem 34].
The probabilistic description of the fine topology is given by the following key result, according to [33, Chapter II, Theorem 4.8], or [66, Proposition 10.8 and Exercise 10.18], and [21, Corollary A.10]. The second assertion states the relation between excessive functions and right-continuous supermartingales (cf. e.g., [18, Proposition 1], see also [19]).
@@
A remarkable property which has been relatively overlooked in the literature is that given a right process $X$ on a general Lusin measurable space $(E,\mathcal{B})$, the operators $P_{t}$ have the property that $P_{t}f$ is finely continuous whenever $f$ is finely continuous, bounded, and $\mathcal{B}^{u}$-measurable; see e.g. [66, Exercise 10.24], [33, Exercise 4.14], or [19, Lemma 2.8] for a proof of this result. In other words, $P_{t}$ is Feller with respect to the fine topology.
@@
Next, by [41, Chapter IV, Theorem 34; see also pages 91-92], we have that $W^{E}$ is a universally measurable subset of ${E}^{D}$. Consequently, we have
@@ 4. $\circ$
The following two results generalize several classical tools due to Dynkin [43, Chapter XII, Sections 4 5] that concern the balayage operator and are key ingredients in solving the stochastic parabolic Dirichlet problem in Section 3.2.
@@
Let $A\in\mathcal{B}$ and set $f:=B_{\alpha}^{A}u$. In order to show that $f$ is finely continuous (resp. on $({E\setminus A})^{\circ_{f}}$) it is sufficient to prove that if $\varepsilon>0$ then $x$ is irregular for $V:=f^{-1}([f(x)+\varepsilon,\infty))$ and for $W:=f^{-1}((-\infty,f(x)-\varepsilon])$, for every $x\in E$ (resp. $x\in({E\setminus A})^{\circ_{f}}$). We treat only the case of $V$, the one for $W$ being similar. Fix $x\in E$. Let $(V_{n})_{n}$ be an increasing sequence of finely closed sets such that $T_{V_{n}}\searrow T_{V}\;\mathbb{P}_{x}$-a.s., which exists by e.g. [66, Chapter I, Exercise (10.30)] [33, Theorem 10.19]. By the zero-one law ([66, Chapter I, Corollary (3.11)] or [33, Proposition 5.17]), $x$ is either regular for $V$, i.e. $\mathbb{P}_{x}([T_{V}=0])=1$, or irregular for $V$, i.e. $\mathbb{P}_{x}([T_{V}=0])=0$.
@@
and also that $T_{V_{n}}+T_{A}\circ\theta_{T_{V_{n}}}\searrow_{n}T_{A}$ $\mathbb{P}_{x}$-a.s., see e.g. [66, p. 65] or [33, (10.3)].
@@
Let us now prove (i). To this end, let $x\in({E\setminus A})^{\circ_{f}}$, so that by 6.20 we have $\mathbb{P}_{x}(T_{A}>0)=1$. Using again that $T_{A}$ is a terminal time ([66, p. 65]) and (6.3), we get
@@
The fact that $M$ is càdlàg $\mathbb{P}_{x}$-a.s. now follows from a standard result of convergence for increasing families of càdlàg supermartingales, see e.g. [41, Chapter VI, Theorem 18] or [42, Part II, Chapter IV, Section 4].
@@
    The map $E\ni x\mapsto\delta_{x}\circ U_{\beta}\in Exc(\mathcal{U}_{\beta})$ is a measurable embedding of $E$ into $E_{1}$ and every $\mathcal{U}_{\beta}$-excessive function $v$ has an extension $v_{1}$ to $E_{1}$, defined as $v_{1}(\xi):=L^{\beta}(\xi,v)$. The set $E_{1}$ is endowed with the $\sigma$-algebra $\mathcal{B}_{1}$ generated by the family $\{v_{1}:\;v\in\mathcal{E}(\mathcal{U}_{\beta})\}$. In addition, as in [16, Sections 1.1 and 1.2], there exists a unique resolvent of kernels $\mathcal{U}^{1}=(U^{1}_{\alpha})_{\alpha>0}$ on $(E_{1},\mathcal{B}_{1})$ which is an extension of $\mathcal{U}$ in the sense that
@@
The existence of a right process on a larger space, more precisely on $E_{1}$, is given by the following result, for which we refer to [29, (2.3)], in [13, Sections 1.7 and 1.8], [16, Theorem 1.3], and [17, Section 3]; see also [69].
@@ 6.2 Choquet capacity
is a Choquet capacity on $(E,\mathcal{T})$. Moreover, if $X$ has left limits in $(E,\mathcal{T})$ $\mathbb{P}_{\nu}$-a.s., then the capacity ${\sf Cap}^{\alpha}_{\nu}$ is tight.
For convenience, let us briefly recall the notion of capacity (for details see e.g. [41, Ch. III] and [13, Appendix A.1]):
@@ 6.3.2 Restriction, trivial extension, and trivial modification of the process
We denote by $\mathcal{S}_{\alpha}$ the set of all strongly supermedian functions on $E$ (with respect to $\mathcal{U}_{\alpha}$). Further, we present several basic properties of the strongly supermedian functions, cf. e.g. [12] and [13], and [26].
@@
For the construction of the process $X^{\prime}$, the extension of $X$ from $F$ to $E$, we refer to [57, p. 118].
@@ 6.5 Adding jumps by perturbation with kernels
The next proposition is a version of Proposition 4.5 from [27].
@@
The proof now finishes by noticing that $X^{K}$ corresponds to the process constructed in [50, Theorem 2.5], hence it is in fact a Hunt process.
@@
(i). First of all, we can apply [44, Chapter III, 1.3] two times in a row, first for the perturbation $L_{1}-c$ and then for the perturbation $L_{1}-c+K$, to ensure that $\left(L_{1}-c+K,\;D(L_{1})\right)$ generates a $C_{0}$-semigroup on $L^{1}(\mu)$, which we denote by $(S_{t})_{t\geq 0}$. Furthermore, by [44, Chapter III, 1.7] we have
@@ Acknowledgements.
We gratefully acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.

### 11 Mar 2026 v2

@@ 2. $(T4^{\prime})$
Further, $\mathbf{H_{\sf\lesssim u}}$ follows from [64, Theorem 33] presented above, and 4.6, (i).
@@ 6.1 Fundamentals on right processes and their potential theory
Let $(E,\mathcal{B})$ be a Lusin measurable space, i.e. it is measurably isomorphic to a Borel subset of a compact metric space endowed with the corresponding Borel $\sigma$-algebra. We denote by $(b)p\mathcal{B}$ the set of all numerical, (bounded) positive $\mathcal{B}$-measurable functions on $E$. Throughout, by $\mathcal{U}=(U_{\alpha})_{\alpha>0}$ we denote a resolvent family of (sub-)Markovian kernels on $(E,\mathcal{B})$; see e.g. [14, Subsection 1.1]. If $\beta>0$, we set $\mathcal{U}_{\beta}:=(U_{\beta+\alpha})_{\alpha>0}$.
@@
    Let $(E,\mathcal{B})$ be a Lusin measurable space, and $(f_{n})_{n\geq 1}\subset\mathcal{B}$ be such that it separates the points of $E$. Then the topology on $E$ generated by $(f_{n})_{n\geq 1}$ is a Lusin topology; see e.g. [60, Theorem A2.12] for the similar Radonian case.
@@
Indeed, one can construct such $\tilde{X}$ by considering first a second process equivalent to $X$ which is of function space type, according to [33, Definition 4.2 and Theorem 4.3], and then simply deleting from the path space of this second process the set $\Lambda$ of those paths that are not right-continuous with respect to $\tau$. Such a removal is allowed since $\Lambda$ is universally measurable according to [40, Ch. IV, Theorem 34].
The probabilistic description of the fine topology is given by the following key result, according to [33, Chapter II, Theorem 4.8], or [60, Proposition 10.8 and Exercise 10.18], and [22, Corollary A.10]. The second assertion states the relation between excessive functions and right-continuous supermartingales (cf. e.g., [19, Proposition 1], see also [20]).
@@
A remarkable property which has been relatively overlooked in the literature is that given a right process $X$ on a general Lusin measurable space $(E,\mathcal{B})$, the operators $P_{t}$ have the property that $P_{t}f$ is finely continuous whenever $f$ is finely continuous, bounded, and $\mathcal{B}^{u}$-measurable; see e.g. [60, Exercise 10.24], [33, Exercise 4.14], or [20, Lemma 2.8] for a proof of this result. In other words, $P_{t}$ is Feller with respect to the fine topology.
@@
Next, by [40, Chapter IV, Theorem 34; see also pages 91-92], we have that $W^{E}$ is a universally measurable subset of ${E}^{D}$. Consequently, we have
@@ 4. $\circ$
The following two results generalize several classical tools due to Dynkin [42, Chapter XII, Sections 4 5] that concern the balayage operator and are key ingredients in solving the stochastic parabolic Dirichlet problem in Section 3.2.
@@
Let $A\in\mathcal{B}$ and set $f:=B_{\alpha}^{A}u$. In order to show that $f$ is finely continuous (resp. on $({E\setminus A})^{\circ_{f}}$) it is sufficient to prove that if $\varepsilon>0$ then $x$ is irregular for $V:=f^{-1}([f(x)+\varepsilon,\infty))$ and for $W:=f^{-1}((-\infty,f(x)-\varepsilon])$, for every $x\in E$ (resp. $x\in({E\setminus A})^{\circ_{f}}$). We treat only the case of $V$, the one for $W$ being similar. Fix $x\in E$. Let $(V_{n})_{n}$ be an increasing sequence of finely closed sets such that $T_{V_{n}}\searrow T_{V}\;\mathbb{P}_{x}$-a.s., which exists by e.g. [60, Chapter I, Exercise (10.30)] [33, Theorem 10.19]. By the zero-one law ([60, Chapter I, Corollary (3.11)] or [33, Proposition 5.17]), $x$ is either regular for $V$, i.e. $\mathbb{P}_{x}([T_{V}=0])=1$, or irregular for $V$, i.e. $\mathbb{P}_{x}([T_{V}=0])=0$.
@@
and also that $T_{V_{n}}+T_{A}\circ\theta_{T_{V_{n}}}\searrow_{n}T_{A}$ $\mathbb{P}_{x}$-a.s., see e.g. [60, p. 65] or [33, (10.3)].
@@
Let us now prove (i). To this end, let $x\in({E\setminus A})^{\circ_{f}}$, so that by 6.20 we have $\mathbb{P}_{x}(T_{A}>0)=1$. Using again that $T_{A}$ is a terminal time ([60, p. 65]) and (6.3), we get
@@
The fact that $M$ is càdlàg $\mathbb{P}_{x}$-a.s. now follows from a standard result of convergence for increasing families of càdlàg supermartingales, see e.g. [40, Chapter VI, Theorem 18] or [41, Part II, Chapter IV, Section 4].
@@
    The map $E\ni x\mapsto\delta_{x}\circ U_{\beta}\in Exc(\mathcal{U}_{\beta})$ is a measurable embedding of $E$ into $E_{1}$ and every $\mathcal{U}_{\beta}$-excessive function $v$ has an extension $v_{1}$ to $E_{1}$, defined as $v_{1}(\xi):=L^{\beta}(\xi,v)$. The set $E_{1}$ is endowed with the $\sigma$-algebra $\mathcal{B}_{1}$ generated by the family $\{v_{1}:\;v\in\mathcal{E}(\mathcal{U}_{\beta})\}$. In addition, as in [17, Sections 1.1 and 1.2], there exists a unique resolvent of kernels $\mathcal{U}^{1}=(U^{1}_{\alpha})_{\alpha>0}$ on $(E_{1},\mathcal{B}_{1})$ which is an extension of $\mathcal{U}$ in the sense that
@@
The existence of a right process on a larger space, more precisely on $E_{1}$, is given by the following result, for which we refer to [29, (2.3)], in [14, Sections 1.7 and 1.8], [17, Theorem 1.3], and [18, Section 3]; see also [63].
@@ 6.2 Choquet capacity
is a Choquet capacity on $(E,\mathcal{T})$. Moreover, if $X$ has left limits in $(E,\mathcal{T})$ $\mathbb{P}_{\nu}$-a.s., then the capacity ${\sf Cap}^{\alpha}_{\nu}$ is tight; cf. [52], [16], and [30].
For convenience, let us briefly recall the notion of capacity (for details see e.g. [40, Ch. III] and [14, Appendix A.1]):
@@ 6.3.2 Restriction, trivial extension, and trivial modification of the process
We denote by $\mathcal{S}_{\alpha}$ the set of all strongly supermedian functions on $E$ (with respect to $\mathcal{U}_{\alpha}$). Further, we present several basic properties of the strongly supermedian functions, cf. e.g. [13], [14], [15], and [26].
@@
For the construction of the process $X^{\prime}$, the extension of $X$ from $F$ to $E$, we refer to [53, p. 118].
@@ 6.5 Adding jumps by perturbation with kernels
The next proposition is a version of Proposition 4.5 from [27]; see also Proposition 3.4 from [28].
@@
The proof now finishes by noticing that $X^{K}$ corresponds to the process constructed in [49, Theorem 2.5], hence it is in fact a Hunt process.
@@
(i). First of all, we can apply [43, Chapter III, 1.3] two times in a row, first for the perturbation $L_{1}-c$ and then for the perturbation $L_{1}-c+K$, to ensure that $\left(L_{1}-c+K,\;D(L_{1})\right)$ generates a $C_{0}$-semigroup on $L^{1}(\mu)$, which we denote by $(S_{t})_{t\geq 0}$. Furthermore, by [43, Chapter III, 1.7] we have
@@ Acknowledgements.
We gratefully acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283. I.C. was partially supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS-UEFISCDI, project number PN-IV-P2-2.1-T-TE-2023-1704, within PNCDI IV.